A proton, moving with a velocity of 31, collides elastically with another proton that is initi al rest. If the two protons have equal speeds after the collision. 07:- Find the speed of each proton after the collision (A)0.71ms- (B) 1.4ms1 (C)2. 1ms-1 (D) 2.8ms (E)3.5ms- 08:-The direction of the velc ectors after the collision (A(S5,-35) (B)(30 6 (C(45.-45) (D)(60,-30) (E50-40)
Answer #1

Nomenclature is like this

u1xi = initial speed of moving proton in x-direction = 3

u2xi = initial speed of stationary proton in x-direction = 0

u1yi = initial speed of moving proton in y-direction = 0

u2yi = initial speed of stationary proton in y-direction = 0

V1, V2 = final speed of both protons = V

\theta1 = final direction of initially moving proton

\theta2 = final direction of initially stationary proton

Also see that one proton will go Counterclockwise from +ve axis and other proton will go clockwise from +x-axis, which means  \theta1 will be +ve and \theta2 will be -ve.

In a perfectly elastic collision

Using momentum conservation in x direction,
Pi = Pf
m1*u1xi + m2*u2xi = m1*V1x + m2*V2x
given that m1 = m & m2 = m

u1xi = 3 m/sec & u2xi = 0 (since stationary)

V1x = V*cos \theta1 & V2x = V*cos \theta2
m*u1xi + m*u2xi = m*V*cos \theta1 + m*V*cos \theta2
3 - 0 = V*cos \theta1 + V*cos \theta2
V*cos \theta1 + V*cos \theta2 = 3

Using momentum conservation in y direction,
Pi = Pf
m1*u1yi + m2*u2yi = m1*V1y + m2*V2y
given that m1 = m & m2 = m

u1yi = 0 m/sec & u2yi = 0 m/sec

V1y = V*sin \theta1 & V2y = -V*sin \theta2
m*u1i + m*u2i = m*V*sin \theta1 - m*V*sin \theta2

0 - 0 = V*sin \theta1 - V*sin \theta2

V*sin \theta1 - V*sin \theta2 = 0

V*sin \theta1 = V*sin \theta2

sin \theta1 = sin \theta2

\theta1 = \theta2 = \theta

from given option we can see that

\theta = 45deg

(\theta1, \theta2) = (45, -45)

Correct answer of Q8. is option C.

Now Using above relation in 1st equation

V*cos \theta1 + V*cos \theta2 = 3

V*cos \theta + V*cos \theta = 3

2*V*cos \theta = 3

V = 3/(2*cos 45 deg)

V = 2.1 m/sec

Correct option of Q7 is option C

Please Upvote.

Comment below if you have any query.